# Learning-based Axial Motion Magnification

### v1

@@ Abstract
Video motion magnification amplifies invisible small motions to be perceptible, which provides humans with spatially dense and holistic understanding about small motions from the scene of interest. This is based on the premise that magnifying small motions enhances the legibility of the motion. In the real world, however, vibrating objects often possess complex systems, having complex natural frequencies, modes, and directions. Existing motion magnification often fails to improve the legibility since the intricate motions still retain complex characteristics even when magnified, which distracts us from analyzing them. In this work, we focus on improving the legibility by proposing a new concept, axial motion magnification, which magnifies decomposed motions along the user-specified direction. Axial motion magnification can be applied to various applications where motions of specific axes are critical, by providing simplified and easily readable motion information. We propose a novel learning-based axial motion magnification method with the Motion Separation Module that enables to disentangle and magnify the motion representation along axes of interest. Further, we build a new synthetic training dataset for the axial motion magnification task. Our proposed method improves the legibility of resulting motions along certain axes, while adding additional user controllability. Our method can be directly adopted to the generic motion magnification and achieves favorable performance against competing methods. Our project page is available at https://axial-momag.github.io/axial-momag/.
@@ 1 Introduction
Motions are always present in our surroundings. Among them, small motions often convey important signals in practical applications, e.g., building structure health monitoring [4, 5, 8, 6, 7, 25], machinery fault detection [26, 36, 30], sound recovery [9], and healthcare [1, 2, 22, 11, 14]. Video motion magnification [19, 38, 23, 37, 40] is the technique to amplify subtle motions in a video, revealing details of motion that are hard to perceive with the naked eyes. This allows users to grasp spatially dense and holistic behavior information of the scene of interest instantly, as long as the resulting motion is simple and easily interpretable. However, in practice, vibrating objects in the real world often possess complex systems, having complex natural frequencies, modes, and directions [24]. Even when magnified, the intricate movement within videos persists, which restricts the advantages of motion magnification because the key underlying premise of its effectiveness is based on the legibility of the magnified motion in aforementioned applications, i.e., effectively understanding the way objects move.
@@
Figure 1: Importance of axial motion magnification. When identifying faults in rotating machinery, analysis of the vulnerable axial vibration is critical [21]. Existing motion magnification methods, e.g., Deep Motion Magnification (DMM) [23] amplifies motions along all axes, which yield artifacts driven by amplifying the dominant motion in this sample. It hinders the analyses of vulnerable axial vibration. This motivates the importance of our axial motion magnification that magnifies decomposed motions along a user-specified axis. We magnify the axial vibration only, achieving artifact-free magnified results and the legibility of critical motions. For the visualization purpose, we overlay the sampled trajectories obtained from the Kanade-Lucas-Tomasi (KLT) Tracker [20].
@@
In this work, we focus on improving the legibility of magnified motion by proposing a novel concept, axial motion magnification, which magnifies decomposed motions along a user-specified direction. All the existing works, e.g.,  [40, 37, 38, 23, 16, 29], have overlooked this key importance of the legibility according to axes in practice. There are many practical cases where the importance of motion varies according to axes. For example, in the fault detection application of machines in Fig. 1, even small motions along the vulnerable axis are critical while bigger and dominant rotational motions are not [21]. Likewise, many apparatus consisting of natural or artificial materials often have vulnerable axes due to the asymmetry property of microstructures, e.g., fracture toughness [35, 3, 17]. This motivates us to separately analyze motions according to axes.
@@
Specifically, we propose a novel learning-based axial motion magnification method, where the motions in a user-specified axis are magnified. Our method can independently magnify small motions along two orthogonal orientation axes with two independent magnification factors for each axis, which facilitates the analysis of complex small motions in the lens of axes favorable to the user. To this end, we propose Motion Separation Module (MSM) that disentangles the motion representation into two orthogonal orientations and manipulates it into the direction specified by the user. Additionally, for training the neural network, we develop and build a new synthetic dataset for the axial motion magnification task. Thereby, our proposed approach adds user control capability, while improving the legibility of resulting motions along a certain axis. We can directly adopt our method to the generic motion magnification task and achieve favorable performance against competing methods. We summarize our main contributions as follows:
@@
  • We propose the first learning-based axial motion magnification method, which allows us to selectively amplify small motions along a specific direction.
  • We construct a new synthetic dataset to train the new axial motion magnification model.
@@
  • We conduct analyses and find that adopting the Motion Separation Module (MSM) is effective in distinguishing small motions from noise.
@@ 2 Related Work
Wu et al. [40] re-formulate the motion magnification task as an Eulerian method that represents motion by intensity changes of pixels at each fixed location without actual movement [12]. The Eulerian approach, e.g., [40, 37, 38, 41, 31, 23, 32, 33, 34], becomes standard in motion magnification due to its noise robustness, sensitivity to small motions, and simplified system by avoiding the challenges of the warp and inpaint approach for filling holes and handling occlusions. The system of the Eulerian methods typically consists of motion representation, manipulation, and reconstruction. The previous works can be categorized into two main focuses: 1) proposing motion representations or 2) motion manipulation methods.
@@
To deal with it, Oh et al[23] first coined learning-based video motion magnification, called Deep Motion Magnification (DMM), by modeling motion representation with deep neural networks. Since no real data exists for training video motion magnification, they propose a method to build motion magnification synthetic data. With the development, other learning-based variants [16, 29] have been proposed by focusing on neural network architectures. These approaches demonstrate promising results by effectively handling diverse challenging scenarios such as occlusion and noisy inputs. Also, the motion magnification factors of the data-driven approaches can be controlled by the way the synthetic dataset is generated, while those of the traditional methods [40, 37] are theoretically restricted.
Figure 2: Proposed network architecture. (a) The Encoder outputs features from input images and the features are fed to the Texture branch and Motion Separation Module (MSM). (b) Using weight-shared 1D convolutions, the Shape branch extracts shape representations along the $x$ and $y$-axis, which are sequentially projected by the user-defined direction $\phi$ and ϕsubscriptitalic-$\phi_{\perp}$, respectively. The Decoder predicts the axially magnified image from the outputs from the Texture branch and MSM.
@@
In this work, we present a new notion of motion magnification by disentangling motion axes of the user’s interest. We design a neural architecture to induce disentanglement of motion in oriented axes. Also, to train such a model, we propose the synthetic data generation pipeline for the axial motion magnification task. In contrast to all the existing Eulerian methods, which overlook the directional legibility of the resulting magnified motions, we add an additional feature to motion magnification.
@@ 3 Learning-based Axial Motion Magnification
In this section, we first briefly discuss preliminaries about generic motion magnification, which refers to the methods that amplify the motion regardless of motion direction, including the prior arts [40, 23, 29, 16] (Sec. 3.1). Then, we re-frame the motion magnification problem in the view of axial motion magnification (Sec. 3.2), and elaborate on our network architecture, synthetic data generation method, and objective functions (Sec. 3.3).
@@ 3.1 Preliminary – Generic Motion Magnification
Following the convention [40, 37], for simplicity, we consider the 1D image intensity being shifted by the displacement function $\delta(x,t)$ parameterized by position $x$ and time $t$. It can be generalized to local translational motion in 2D image [40]. Given an underlying intensity profile function $f(\cdot)$, the 1D image intensity $I(x,t)$ can be represented as
@@
The goal of motion magnification is to synthesize the magnified image intensity $\hat{I}(x,t)$:
@@
where $\alpha$ denotes the magnification factor. The key factor of motion magnification methods lies in the extraction of the displacement function $\delta(x,t)$ from Eq. (1). If $\delta(x,t)$ can be decomposed, we can approximate $\hat{I}(x,t)$ by multiplying $\delta(x,t)$ with the magnification factor $\alpha$ and applying the reverse of the decomposition process. However, it is challenging and ill-posed to extract exact displacements from the observed intensity images [40]. Instead, the prior arts approximately decompose $\delta(x,t)$; for example, Wu et al[40] use the first-order Taylor expansion:
@@ 3.2 Axial Motion Magnification
Figure 3: Synthetic data generation pipeline for axial motion magnification. From the sampled background and foregrounds, each with their own segmentation masks, we compose the previous layer images $\{{\mathbf{L}}^{k}_{1}\}_{k=1}^{K}$ and masks $\{\boldsymbol{\Omega}^{k}_{1}\}_{k=1}^{K}$. To generate next layer images $\{{\mathbf{L}}^{k}_{2}\}_{k=1}^{K}$ and masks $\{\boldsymbol{\Omega}^{k}_{2}\}_{k=1}^{K}$, we apply the random translations to $\{{\mathbf{L}}^{k}_{1}\}_{k=1}^{K}$ and $\{\boldsymbol{\Omega}^{k}_{1}\}_{k=1}^{K}$. Axially magnified layer images $\{\hat{{\mathbf{L}}}^{\phi,k}\}_{k=1}^{K}$ and masks $\{\hat{\boldsymbol{\Omega}}^{\phi,k}\}_{k=1}^{K}$ are also synthesized by translations but with the axially magnified translation parameters. The previous, next, and axially magnified layer images and masks are then superimposed into a single image to yield $\textbf{I}_{1}$, $\textbf{I}_{2}$, and $\hat{\textbf{I}}^{\phi}$, respectively. The dataset also include angles $\phi$ and the object-wise magnification maps $\boldsymbol{\Lambda}$ generated by superimposing $\{\boldsymbol{\alpha}^{k}\}_{k=1}^{K}$ with $\{\boldsymbol{\Omega}_{1}^{k}\}_{k=1}^{K}$.
@@
where $\alpha^{\phi}\geq 0$ denotes the axial magnification factor and $\boldsymbol{\delta}^{\phi}(\textbf{x},t)$ represents the projection of $\boldsymbol{\delta}(\textbf{x},t)$ onto a 2D directional unit vector $\textbf{p}^{\phi}$ of the angle $\phi$, i.e., the motion component. We can break down the motion component $\boldsymbol{\delta}^{\phi}(\textbf{x},t)$ into:
$\boldsymbol{\delta}^{\phi}(\textbf{x},t)=\operatorname{proj}_{\mathbf{\textbf{p}^{\phi}}}{\boldsymbol{\delta}(\textbf{x},t)},$ (5)
which we refer to as angular modulation. Thus, the problem boils down to how to extract $\boldsymbol{\delta}^{\phi}$ from the input video.
@@ 3.3 Neural Networks and Training
Since the previous learning-based methods [23, 29, 16] were designed for generic motion magnification, we design a new neural network architecture and a training dataset to learn the motion representation proportional to $\boldsymbol{\delta}^{\phi}$.
@@
Our whole architecture consists of Encoder, Texture branch, Shape branch, Manipulator, and Decoder similar to DMM [23] (see Fig. 2-(a)), where texture represents color and texture-related information while shape represents scene structure-related information that later leads to motion $\boldsymbol{\delta}$ [23]. To extract axial shape representations, we design Motion Separation Module (MSM) consisting of the completely re-designed and dedicated Shape branch and Manipulator as depicted in Fig. 2-(b). In MSM, instead of extracting a single specified direction $\boldsymbol{\delta}^{\phi}$, we design to extract its orthogonal direction $\boldsymbol{\delta}^{\phi_{\perp}}$ as well.
@@
Given consecutive input images $\textbf{I}_{t}\in\mathbb{R}^{H\times W\times 3}$ at $t=1$ and $t=2$, texture representations $\textbf{T}_{t}\in\mathbb{R}^{H/4\times W/4\times 32}$ are obtained by $\textbf{T}_{t}=F(E(\textbf{I}_{t}))$, where $E(\cdot)$ and $F(\cdot)$ denote the Encoder and the Texture branch, respectively. The outputs of $E$ are fed into MSM.
@@
To extract he motion representations along two orthogonal orientations and manipulate them based on the user-defined angle, we grant the learnable parameters to learn the directionality in MSM. Our Shape branch $G(\cdot)$ first extracts the axial shape representations along the canonical $x$ and $y$-axes by applying weight-shared 1D convolutions but with spatially transposing the convolution kernels, yielding $[\textbf{S}^{x}_{t},\textbf{S}^{y}_{t}]$=$G(E(\textbf{I}_{t}))$ where $\textbf{S}^{x}_{t},\textbf{S}^{y}_{t}\in\mathbb{R}^{H/2\times W/2\times 32}$. Then, these are projected according to the user-defined direction $\phi$ by the Projection layer, which produces axial shape representations, i.e., $\textbf{S}^{\phi}_{t}$ and $\textbf{S}^{\phi_{\perp}}_{t}$.
Figure 4: [Left] Imposing an imbalance on a rotor, [Right] Qualitative results in axial motion magnification scenario. We attach weights to a rotor to impose an imbalance and acquire rotor imbalance sequence, which has axial vibrations. Then, we amplify only the motion of rotor’s axial direction with the magnification factor $\alpha=40$, using ours and phase-based method [37]. We also show the magnified result of DMM [23] as one reference result of generic motion magnification methods. Our method produces magnified output frames without artifacts and exhibits the $x$-t graph showing clearly legible axial vibrations, while the magnified results of phase-based method and DMM both suffer from severe artifacts and have unclear axial vibrations in the $x$-t graph.
@@
The Manipulator $M(\cdot)$ calculates the difference of the axial shape representations and magnifies them by multiplying the axial magnification factor $\alpha^{\phi}$. Then, these manipulated representations are fed into subsequent 1D convolutions, and added to the axial shape representation $\textbf{S}^{\phi}_{2}$. For ϕsubscriptitalic-$\phi_{\perp}$, we use the same manipulator, of which weights are shared but spatially transposed, for applying $\alpha^{\phi_{\perp}}$. Note that, with this separation of $\phi$ and ϕsubscriptitalic-$\phi_{\perp}$, we can set the magnification factors $\alpha^{\phi}$ and $\alpha^{\phi_{\perp}}$ independently, enabling broad applications of controls as another benefit. Finally, the Decoder $D(\cdot)$ predicts the axially magnified output frame $\tilde{\textbf{I}}^{\phi}$ as
$\tilde{\textbf{I}}^{\phi}=D\left(\textbf{T}_{2},M(\textbf{S}^{\phi}_{1},\textbf{S}^{\phi}_{2},\alpha^{\phi}),M(\textbf{S}^{\phi_{\perp}}_{1},\textbf{S}^{\phi_{\perp}}_{2},\alpha^{\phi_{\perp}})\right).$ (6)
Note that the Decoder is designed and trained to deal with both $S^{\phi}$ and $S^{\phi_{\perp}}$ in a unified manner. This enables the network to conduct both generic and axial motion magnification, given the user setting of the angle $\phi$.
@@
In the real world, acquiring consecutive images and magnified images at the same time is impossible. Due to this, DMM [23] proposes a synthetic training dataset for the generic motion magnification task. While the dataset is effective for learning generic motion magnification, it is not effective and sufficient to induce the disentanglement of the axial property we need. Thus, we propose a new synthetic dataset specifically designed for learning the axial motion magnification, where the motion between $\textbf{I}_{1}$ and $\hat{\textbf{I}}^{\phi}$ is associated with the angle $\phi$ and axial magnification factor vector $\boldsymbol{\alpha}{=}(\alpha^{\phi};\alpha^{\phi_{\perp}})$. Motivated by the established synthetic dataset generation protocol of DMM, we synthesize the training data pairs using the widely adopted simple copy-paste method [23, 13].

### v3

# Learning-based Axial Video Motion Magnification

@@ Abstract
Video motion magnification amplifies invisible small motions to be perceptible, which provides humans with a spatially dense and holistic understanding of small motions in the scene of interest. This is based on the premise that magnifying small motions enhances the legibility of motions. In the real world, however, vibrating objects often possess convoluted systems that have complex natural frequencies, modes, and directions. Existing motion magnification often fails to improve legibility since the intricate motions still retain complex characteristics even after being magnified, which may distract us from analyzing them. In this work, we focus on improving legibility by proposing a new concept, axial motion magnification, which magnifies decomposed motions along the user-specified direction. Axial motion magnification can be applied to various applications where motions of specific axes are critical, by providing simplified and easily readable motion information. To achieve this, we propose a novel Motion Separation Module that enables to disentangle and magnify the motion representation along axes of interest. Furthermore, we build a new synthetic training dataset for the axial motion magnification task. Our proposed method improves the legibility of resulting motions along certain axes by adding a new feature: user controllability. Axial motion magnification is a more generalized concept; thus, our method can be directly adapted to the generic motion magnification and achieves favorable performance against competing methods. Our project page is available at https://axial-momag.github.io/axial-momag/.
@@ 1 Introduction
Motions are always present in our surroundings. Among them, small motions often convey important signals in practical applications, e.g., building structure health monitoring [4, 5, 8, 6, 7, 27], machinery fault detection [28, 38, 32], sound recovery [9], and healthcare [1, 2, 23, 11, 16]. Video motion magnification [20, 40, 24, 39, 42] is the technique to amplify subtle motions in a video, revealing details of motion that are hard to perceive with the naked eyes. This allows users to grasp spatially dense and holistic behavior information of the scene of interest instantly, as long as the resulting motion is simple and easily interpretable. However, in practice, vibrating objects in the real world often possess complex systems that have complex natural frequencies, modes, and directions [25]. Even after being magnified, the intricate movement within a video persists, which restricts the advantages of motion magnification because the key underlying premise of its effectiveness is based on the legibility of the magnified motion in aforementioned applications, i.e., effectively understanding the way objects move.
@@
Figure 1: Importance of axial motion magnification. When identifying faults in rotating machinery, analysis of the vulnerable axial vibration is critical [22]. Existing learning-based methods [24, 17, 26, 31] amplify motions along all axes, which yield artifacts. It hinders the analyses of vulnerable axial vibration. This motivates the importance of our axial motion magnification that magnifies decomposed motions along a user-specified axis. We magnify the axial vibration only, achieving artifact-free results and the legibility of critical motions. For the visualization purpose, we overlay the sample trajectories obtained from the Kanade-Lucas-Tomasi (KLT) Tracker [21].
@@
In this work, we focus on improving the legibility of magnified motion by proposing a novel concept, axial motion magnification, which magnifies decomposed motions along the user-specified direction. All the existing works, e.g.,  [42, 39, 40, 24, 17, 31, 26], have overlooked this key importance of the legibility according to axes in practice. There are many practical cases where the importance of motion varies according to axes. For example, in the fault detection application of machines in Fig. 1, even small motions along the vulnerable axis are critical while bigger and dominant rotational motions are not [22]. Likewise, many apparatus consisting of natural or artificial materials often have vulnerable axes due to the asymmetry property of microstructures, e.g., fracture toughness [37, 3, 18]. This motivates us to separately analyze motions according to axes.
@@
Specifically, we propose a novel learning-based axial motion magnification method, where the motions in a user-specified axis are magnified. Our method can independently magnify small motions along two orthogonal orientation axes with two independent magnification factors for each axis, which facilitates the analysis of complex small motions in the lens of axes favorable to the user. To this end, we propose the Motion Separation Module (MSM) that disentangles the motion representation into two orthogonal orientations and manipulates it into the direction specified by the user. For training the proposed neural network, we develop and build a new synthetic dataset for the axial motion magnification task. Thereby, our proposed approach adds a new user control feature, which improves the legibility of resulting motions along a certain axis. This allows our axial motion magnification becomes a generalization of the existing generic motion magnification. Thus, our method can be directly adopted to the generic motion magnification task and achieve favorable performance against competing methods. We summarize our main contributions as follows:
@@
  • We propose the new concept, learning-based axial motion magnification, which allows us to selectively amplify small motions along a specific direction.
@@
  • We propose and analyze the Motion Separation Module (MSM) for the axial motion magnification. We find that adopting the MSM is effective in not only axial magnification but also distinguishing small motions from noise.
  • We propose the way to synthesize a new synthetic dataset to train the new axial motion magnification model.
@@ 2 Related Work
Wu et al. [42] re-formulate the motion magnification task as an Eulerian method that represents motion by intensity changes of pixels at each fixed location without actual movement [12]. The Eulerian approach, e.g., [42, 39, 40, 43, 33, 24, 34, 35, 36], becomes standard in motion magnification due to its noise robustness, sensitivity to small motions, and simple system by avoiding challenging warp and inpaint approach for filling holes and handling occlusions. The system of the Eulerian methods typically consists of motion representation, manipulation, and reconstruction. The previous works can be categorized into two main focuses: 1) proposing motion representations or 2) motion manipulation methods.
@@
To deal with, Oh et al[24] first coined learning-based video motion magnification, called Deep Motion Magnification (DMM), by modeling motion representation with deep neural networks. As no real data exists for training video motion magnification, they propose a method to build motion magnification synthetic data. With the development, other learning-based variants [17, 31] have been proposed, focusing on neural network architectures. These approaches demonstrate promising results by effectively handling diverse challenging scenarios such as occlusion and noisy inputs. Also, the motion magnification factors of the data-driven approaches can be controlled by the way the synthetic dataset is generated, while those of the traditional methods [42, 39] are theoretically restricted.
@@
In this work, we present a new notion of motion magnification by disentangling motion axes of the user’s interest. We design a neural architecture to induce disentanglement of motion in oriented axes. Also, to train such a model, we propose the synthetic data generation pipeline for the axial motion magnification task. In contrast to all the existing Eulerian methods, which overlook the directional legibility of the resulting magnified motions, we add a novel feature to motion magnification.
@@ 3 Learning-based Axial Motion Magnification
In this section, we first briefly discuss preliminaries about generic motion magnification, which refers to the methods that amplify the motion regardless of motion direction, including the prior arts [42, 24, 31, 17, 26] (Sec. 3.1). Then, we re-frame the motion magnification problem in the view of axial motion magnification (Sec. 3.2), and elaborate on our network architecture, and synthetic data generation method (Sec. 3.3).
@@ 3.1 Preliminary – Generic Motion Magnification
Following the convention [42, 39], for simplicity, we consider the 1D image intensity being shifted by the displacement function $\delta(x,t)$ which is parameterized by position $x$ and time $t$. It can be generalized to local translational motion in 2D image [42]. Given an underlying intensity profile function $f(\cdot)$, the 1D image intensity $I(x,t)$ can be represented as
@@
The goal of motion magnification is to synthesize the magnified image $\hat{I}(x,t)$:
@@
where $\alpha$ denotes the magnification factor. The key factor of motion magnification methods lies in the extraction of the displacement function $\delta(x,t)$ from Eq. (1). If $\delta(x,t)$ can be decomposed, we can approximate $\hat{I}(x,t)$ by multiplying $\delta(x,t)$ with the magnification factor $\alpha$ and applying the reverse of the decomposition process. However, it is ill-posed problem to extract exact displacements from the observed intensity images [42]. Instead, the prior arts approximately decompose $\delta(x,t)$; for example, Wu et al[42] use the first-order Taylor expansion as:
@@ 3.2 Axial Motion Magnification
where $\alpha^{\phi}\geq 0$ denotes the axial magnification factor and $\boldsymbol{\delta}^{\phi}(\textbf{x},t)$ the projection of $\boldsymbol{\delta}(\textbf{x},t)$ onto a 2D directional unit vector $\textbf{p}^{\phi}$ with the angle $\phi$, i.e., the motion component. We can break down the motion component $\boldsymbol{\delta}^{\phi}(\textbf{x},t)$ into:
$\boldsymbol{\delta}^{\phi}(\textbf{x},t)=\operatorname{proj}_{\mathbf{\textbf{p}^{\phi}}}{\boldsymbol{\delta}(\textbf{x},t)}.$ (5)
Relationship with Generic Motion Magnification
If we obtain $\boldsymbol{\delta}(\textbf{x},t)$, we can determine $\boldsymbol{\delta}^{\phi}(\textbf{x},t)$ and $\boldsymbol{\delta}^{\phi_{\perp}}(\textbf{x},t)$ through the projections onto $\textbf{p}^{\phi}$ and $\textbf{p}^{\phi_{\perp}}$. In this case, we can extend Eq. 4 to represent not only the displacement vector of an angle $\boldsymbol{\delta}^{\phi}(\textbf{x},t)$ but also of its orthogonal direction $\boldsymbol{\delta}^{\phi_{\perp}}(\textbf{x},t)$, as
$\hat{I}^{\phi}(\textbf{x},t)=f(\textbf{x}+\alpha^{\phi}\boldsymbol{\delta}^{\phi}(\textbf{x},t)+\alpha^{\phi_{\perp}}\boldsymbol{\delta}^{\phi_{\perp}}(\textbf{x},t)),$ (6)
where $\alpha_{\phi}$, $\alpha_{\phi_{\perp}}\geq 0$ denotes the axial magnification factors corresponding to the $\phi$ and ϕsubscriptitalic-$\phi_{\perp}$ directions, respectively. This formulation encompasses the various motion magnification scenarios, e.g., axial and generic motion magnifications. Setting $\alpha^{\phi_{\perp}}$ to 00 leads to the formulation resulting in axial motion magnification, while setting $\alpha^{\phi}$ equal to $\alpha^{\phi_{\perp}}$ results in generic motion magnification.
@@ 3.3 Neural Networks and Training
Departing from the previous learning-based methods that are confined to generic motion magnification [24, 31, 17, 26], we introduce a novel neural network architecture and a dedicated training dataset designed to learn two angle-aware motion representations proportional to the motion displacement $\boldsymbol{\delta}^{\phi}$ and $\boldsymbol{\delta}^{\phi_{\perp}}$, respectively. These allow our approach to unveil a distinctive feature: the magnification of motion in user-defined directions while retaining the functionality for generic motion magnification.
@@
Figure 2: Proposed architecture. (a) The Encoder outputs features from input images and the features are fed to the Texture branch and Motion Separation Module (MSM). (b) Using weight-shared 1D convolutions, the Shape branch extracts shape representations along the $x$ and $y$-axes. These representations are fed to the projection layer $P^{\phi}$, which generates axial shape representations, i.e., $\textbf{S}^{\phi}_{t}$ and $\textbf{S}^{\phi_{\perp}}_{t}$. (c) the Manipulator amplifies them by the axial magnification factors and the inverse projection layer $P^{-\phi}$ re-project them onto the $x$ and $y$-axes. Finally, the Decoder predicts the axially magnified image from the outputs from both the Texture branch and MSM.
@@
Our whole architecture consists of Encoder, Texture & Shape branches, Manipulator, and Decoder similar to DMM [24] (see Fig. 2-(a)), where texture represents color and texture-related information while shape represents scene structure-related information that later leads to motion $\boldsymbol{\delta}$ [24]. To extract axial shape representations, we design Motion Separation Module (MSM) consisting of the completely re-designed and dedicated Shape branch and Manipulator as depicted in Fig. 2-(b,c). In MSM, instead of extracting a single specified direction’s $\boldsymbol{\delta}^{\phi}$, we design to extract its orthogonal direction’s $\boldsymbol{\delta}^{\phi_{\perp}}$ as well. This design choice is motivated by the extended axial motion magnification equation Eq. 6 and enables conducting various motion magnifications, including both axial and generic motion magnifications.
@@
Given consecutive input video frames $\textbf{I}_{t}\in\mathbb{R}^{H\times W\times 3}$ at $t=1$ and $t=2$ for example, texture representations $\textbf{T}_{t}\in\mathbb{R}^{H/4\times W/4\times 32}$ are obtained by $\textbf{T}_{t}=F(E(\textbf{I}_{t}))$, where $E(\cdot)$ and $F(\cdot)$ denote the Encoder and the Texture branch, respectively. The outputs of $E$ are fed into MSM. The same output from $E$ is fed into the Texture branch and MSM, respectively.
@@
To extract the motion representations along two orthogonal orientations and manipulate them based on the user-defined angle, we grant the learnable parameters to learn the directionality in MSM. Our Shape branch $G(\cdot)$ first extracts the axial shape representations along the canonical $x$ and $y$-axes by applying weight-shared 1D convolutions but with spatially transposing the convolution kernels, yielding $[\textbf{S}^{x}_{t},\textbf{S}^{y}_{t}]$=$G(E(\textbf{I}_{t}))$ where $\textbf{S}^{x}_{t},\textbf{S}^{y}_{t}\in\mathbb{R}^{H/2\times W/2\times 32}$. Then, these are projected by the projection layer, which produces axial shape representations of $\phi$ and ϕsubscriptitalic-$\phi_{\perp}$ directions, i.e., $\textbf{S}^{\phi}_{t}$ and $\textbf{S}^{\phi_{\perp}}_{t}$. Motivated by the steerable filters [13], where an arbitrarily rotated representation can be synthesized by a linear combination of directional representations, we design the projection layer $P^{\phi}$ with a linear matrix as
$P^{\phi}\left(\left[\begin{array}[]{l}S_{t}^{x}\\ S_{t}^{y}\end{array}\right]\right)=\left[\begin{array}[]{cc}\cos\phi&\sin\phi\\ -\sin\phi&\cos\phi\end{array}\right]\left[\begin{array}[]{l}S_{t}^{x}\\ S_{t}^{y}\end{array}\right]=\left[\begin{array}[]{l}S_{t}^{\phi}\\ S_{t}^{\phi_{\perp}}\end{array}\right].$ (7)
@@
The Manipulator $M(\cdot)$ computes the difference of the axial shape representations and magnifies them by multiplying the axial magnification factors $\alpha^{\phi}$. Then, these manipulated representations are fed into subsequent 1D convolutions, and added to the axial shape representation $\textbf{S}^{\phi}_{2}$. For ϕsubscriptitalic-$\phi_{\perp}$, we use the same manipulator, of which weights are shared but spatially transposed, for applying $\alpha^{\phi_{\perp}}$. Note that, with this separation of $\phi$ and ϕsubscriptitalic-$\phi_{\perp}$, we can set the magnification factors $\alpha^{\phi}$ and $\alpha^{\phi_{\perp}}$ independently, enabling broad applications of controls as another benefit. For the outputs of the Manipulator $\Delta^{\phi},\Delta^{\phi_{\perp}}$, where $\Delta^{\phi}=M(\textbf{S}^{\phi}_{1},\textbf{S}^{\phi}_{2},\alpha^{\phi})$, we re-project them onto the canonical $x$ and $y$-axes by inverse projection layer $P^{-\phi}$, obtaining $\Delta^{x},\Delta^{y}$. Finally, the Decoder $D(\cdot)$ predicts the axially magnified output frame $\tilde{\textbf{I}}^{\phi}$ as
$\tilde{\textbf{I}}^{\phi}=D\left(\textbf{T}_{2},\Delta^{x},\Delta^{y}\right).$ (8)
This network architecture enables the network to conduct both generic and axial motion magnification, given the user setting of the angle $\phi$. The model is trained with the loss function suggested by DMM [24] with a slight modification to impose the loss separately to the x-axis and y-axis shape representations. Details of the loss function can be found in the supplementary material.
Figure 3: Synthetic data generation pipeline for axial motion magnification. From the sampled background and foregrounds, each with their own segmentation masks, we compose the previous layer images $\{{\mathbf{L}}^{k}_{1}\}_{k=1}^{K}$ and masks $\{\boldsymbol{\Omega}^{k}_{1}\}_{k=1}^{K}$. To generate next layer images $\{{\mathbf{L}}^{k}_{2}\}_{k=1}^{K}$ and masks $\{\boldsymbol{\Omega}^{k}_{2}\}_{k=1}^{K}$, we apply the random translations to $\{{\mathbf{L}}^{k}_{1}\}_{k=1}^{K}$ and $\{\boldsymbol{\Omega}^{k}_{1}\}_{k=1}^{K}$. Axially magnified layer images $\{\hat{{\mathbf{L}}}^{\phi,k}\}_{k=1}^{K}$ and masks $\{\hat{\boldsymbol{\Omega}}^{\phi,k}\}_{k=1}^{K}$ are also synthesized by translations but with the axially magnified translation parameters. These images and masks are then superimposed into a single image to yield $\textbf{I}_{1}$, $\textbf{I}_{2}$, and $\hat{\textbf{I}}^{\phi}$, respectively. The dataset also include angles $\phi$ and the object-wise magnification maps $\boldsymbol{\Lambda}$ generated by superimposing $\{\boldsymbol{\alpha}^{k}\}_{k=1}^{K}$ with $\{\boldsymbol{\Omega}_{1}^{k}\}_{k=1}^{K}$.
@@
In the real world, acquiring consecutive images and magnified images at the same time is impossible. Due to this, DMM [24] proposes a synthetic training dataset for the generic motion magnification task. However, this dataset is not sufficient to induce the disentanglement of the axial property we need. Thus, we propose a new synthetic dataset specifically designed for the axial motion magnification, where the motion between $\textbf{I}_{1}$ and $\hat{\textbf{I}}^{\phi}$ is associated with the angle $\phi$ and axial magnification factor vector $\boldsymbol{\alpha}{=}(\alpha^{\phi};\alpha^{\phi_{\perp}})$. Motivated by the synthetic dataset generation protocol of DMM, we synthesize the training data pairs using the widely adopted simple copy-paste method [24, 14].